Consider a stock with dividends that are expected to grow at 20% per year for four years, after which they are expected to grow at 6% per year, indefinitely. The last dividend paid was $2.00, and k = 11%. Calculate the value of this stock using the multistage growth model.
Answer #1

D1=(2*1.2)=2.4

D2=(2.4*1.2)=2.88

D3=(2.88*1.2)=3.456

D4=(3.456*1.2)=4.1472

Value after year 4=(D4*Growth Rate)/(k-Growth Rate)

=(4.1472*1.06)/(0.11-0.06)

=$87.92064

Hence value of stock=Future dividends and value*Present value of discounting factor(rate%,time period)

=2.4/1.11+2.88/1.11^2+3.456/1.11^3+4.1472/1.11^4+$87.92064/1.11^4

=$67.67(Approx).